# Parameterized Controls and Isomorphic Representations of G-Expectations

### v1

@@ 3.3 True G-Brownian
3.4 The Complete Isomorphism
To establish a full isomorphism, we must introduce time-varying adaptive controls.
3.4.1 Adaptive Control Representation
Definition 3.9 (True G-Brownian Motion via Adaptive Controls).
Let $\Theta\subset\mathbb{R}$ be compact convex. Define:
with quadratic variation:
$d\langle B^{\theta}\rangle_{t}=\theta_{t}^{2}dt$ (3.19)
The genuine G-Brownian motion is characterized by:
$\hat{\mathbb{E}}[\phi(B_{t})]=\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[\phi(B_{t}^{\theta})]$ (3.20)
3.4.2 Algebraic Structure Preservation
The isomorphism preserves all algebraic operations:
Theorem 3.10 (Algebraic Isomorphism).
The mapping $\Phi:X\mapsto\{X^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$ preserves:
  1. 1.
    Addition: $\Phi(X+Y)=\Phi(X)+\Phi(Y)$
  2. 2.
    Scalar Multiplication: $\Phi(\lambda X)=\lambda\Phi(X)$
  3. 3.
    Multiplication: $\Phi(X\cdot Y)=\Phi(X)\cdot\Phi(Y)$
  4. 4.
    Function Composition: $\Phi(\phi(X))=\phi(\Phi(X))$
The preservation follows directly from the pathwise definition:
$(X+Y)^{\theta}$ $=X^{\theta}+Y^{\theta}$
$(\lambda X)^{\theta}$ $=\lambda X^{\theta}$
$(X\cdot Y)^{\theta}$ $=X^{\theta}\cdot Y^{\theta}$
$\phi(X)^{\theta}$ $=\phi(X^{\theta})$
Fourth, we define the multivariate composition in a similar manner. Building upon the correspondence established via G-Brownian motion, we derive a corresponding relationship in the space spanned by these two elements, while ensuring that the mapping constitutes a bijection.
3.4.3 Analytic Structure Preservation
The isomorphism also preserves analytic properties:
Theorem 3.11 (Analytic Isomorphism).
@@ 3.4.3 Analytic Structure Preservation
    Conditional Expectation: $\hat{\mathbb{E}}_{t}[X]=\sup_{\theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}|\mathcal{F}_{t}]$
@@
3.5 Stochastic Calculus Equivalence
The isomorphism extends to differential operations:
@@ 3.5.1 Itô Calculus Preservation
Theorem 3.12 (Itô Calculus Isomorphism).
@@
    $\Phi(dB_{t})=\{\theta_{t}dW_{t}\}$
@@
3.5.2 Itô Calculus Isomorphism (Integral Form)
@@
This confirms the isomorphism preserves the integral structure of Itô calculus in L²-sense. ∎
3.6 Visualizing the Isomorphism: From G-SDE to Parameterized Processes
@@
The isomorphism $\Phi$ establishes a complete correspondence between the G-expectation space and the parameterized classical space:
with the mapping defined as:
$\Phi(X)=\{X^{\theta}\}_{\theta\in\Theta}$
@@ 3.7 Algebraic and Analytic Structure Preservation
The isomorphism preserves all algebraic and analytic operations:
@@
The sublinear expectation operators are related through the isomorphism:
@@
$\hat{\mathbb{E}}[X]$ $\sup\limits_{\theta\in\Theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
$\hat{\mathbb{E}}_{t}[X]$ $\sup\limits_{\theta\in\Theta}\mathbb{E}_{\mathbb{P}}[X^{\theta}\mid\mathcal{F}_{t}]$
@@
    Analytic Continuity: The isomorphism preserves convergence properties:
@@
This isomorphism provides:
  • A computational bridge from nonlinear to linear stochastic calculus
  • A theoretical foundation for model uncertainty quantification
  • An operational framework for solving G-SDEs through parameterized classical SDEs
@@ 4 How to Solve G-Stochastic Differential Equations
$dX_{t}^{\theta}=b(t,X_{t}^{\theta})dt+\sigma(t,X_{t}^{\theta})\theta_{t}dW_{t}$ (4.1)
@@
Under standard Lipschitz and growth conditions on $b$ and $\sigma$, for each $\theta\in\mathcal{A}^{\Theta}$, there exists a unique strong solution $X_{t}^{\theta}$.
@@
$dX_{t}=b(t,X_{t})dt+\sigma(t,X_{t})dB_{t}$ (4.2)
where $B_{t}$ is a G-Brownian motion
Theorem 4.1 (Existence and Uniqueness for G-SDE).
Under bounded controls $\theta_{t}\in\Theta$ and standard Lipschitz conditions:
$|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|$ $\leq K|x-y|$
$|b(t,x)|^{2}+|\sigma(t,x)|^{2}$ $\leq K(1+|x|^{2})$
the G-SDE (4.2) admits a unique solution in the quasi-sure sense:
$X_{t}=\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$
Proof.
The proof follows from:
  1. 1.
    For each fixed $\theta\in\mathcal{A}^{\Theta}$, the classical SDE has a unique strong solution $X_{t}^{\theta}$
  2. 2.
    The boundedness of $\Theta$ ensures uniform estimates:
    $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{s}^{\theta}|^{2}\right]<\infty$
  3. 3.
    The solution family $\{X_{t}^{\theta}\}$ satisfies the consistency conditions for G-expectation
where $\theta_{t}\in\Theta$ is an adapted control process with $\Theta\subset\mathbb{R}$ compact and convex.
4.1 Existence and Uniqueness Theorem
Theorem 4.2 (Classical SDE Solution).
@@ 4.1 Existence and Uniqueness Theorem
Then for each $\theta\in\mathcal{A}^{\Theta}$, there exists a unique strong solution $X_{t}^{\theta}$ to (4.1).
@@
The boundedness of $\theta_{t}$ ensures that the modified diffusion coefficient $\sigma(t,x)\theta_{t}$ satisfies:
$|\sigma(t,x)\theta_{t}-\sigma(t,y)\theta_{t}|$ $\leq M|\sigma(t,x)-\sigma(t,y)|\leq MK|x-y|$
$|\sigma(t,x)\theta_{t}|$ $\leq M|\sigma(t,x)|\leq MK(1+|x|)$
Thus the Lipschitz condition (4.3) and growth condition (4.4) hold for the controlled SDE. Existence and uniqueness follow from standard SDE theory. ∎
4.2 G-SDE Solution via Parameterization
Theorem 4.3 (G-SDE Solution).
@@ 4.2 G-SDE Solution via Parameterization
  1. 1.
    For each $\theta\in\mathcal{A}^{\Theta}$, solve the classical SDE (4.1)
  2. 2.
    The boundedness of $\Theta$ ensures uniform estimates:

### v7

@@ 2.4 Parameterized Control Representation
3 The homomorphism
This chapter is devoted to the rigorous construction of the fundamental homomorphism $\Phi$ between the space of random variables under the G-expectation and the space of parameterized families of classical random variables. The central goal is to demonstrate that this mapping not only exists but also preserves the essential algebraic structure of the spaces involved. The construction proceeds in a canonical manner: we first define $\Phi$ explicitly on a dense subset of cylinder functions, establish its key properties there, and then extend it continuously to the entire space. The preservation of addition, scalar multiplication, and multiplication under this extension is then proven.
3.1 Notation and Preliminaries
Let $(\Omega,\mathcal{F},P)$ be a complete probability space supporting a standard Brownian motion $\{W_{t}\}_{t\geq 0}$, and let $\{\mathcal{F}_{t}\}$ be its natural filtration, augmented to satisfy the usual conditions. Let $\Theta\subset\mathbb{R}$ be a compact convex set representing the volatility uncertainty. The set of admissible controls is defined as:
$\mathcal{A}^{\Theta}:=\left\{\theta:[0,\infty)\times\Omega\to\Theta\mid\theta\text{ is }\{\mathcal{F}_{t}\}\text{-progressively measurable and }E_{P}\left[\int_{0}^{T}|\theta_{s}|^{2}ds\right]<\infty\ \forall T>0\right\}.$
For any $\theta\in\mathcal{A}^{\Theta}$, the controlled process is defined by:
$B_{t}^{\theta}=\int_{0}^{t}\theta_{s}dW_{s}.$
The G-expectation space $L_{G}^{p}(\Omega_{T})$ is the completion of the space of cylinder functions under the norm $\|X\|_{L_{G}^{p}}=\hat{E}[|X|^{p}]^{1/p}$, where $\hat{E}$ denotes the G-expectation operator.
A cylinder function is a random variable of the form:
$\xi=\varphi\left(B_{t_{1}\wedge T},B_{t_{2}\wedge T},\cdots,B_{t_{n}\wedge T}\right),\quad\forall n\in\mathbb{N},t_{1},\ldots,t_{n}\in[0,\infty),\varphi\in C_{l.Lip}(\mathbb{R}^{n})$
where $0\leq t_{1}<t_{2}<\ldots<t_{n}\leq T$ and $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is a local Lipschitz continuous function. The collection of all such cylinder functions is denoted by $\mathcal{C}_{T}$.
Lemma 3.1 (Density of Cylinder Functions).
The set $\mathcal{C}_{T}$ is dense in $L_{G}^{p}(\Omega_{T})$ for any $p\geq 1$.
Proof.
This is a foundational result in the theory of G-expectation. See, for example, [1]
3.2 Construction of the homomorphism $\Phi$
We now construct the mapping $\Phi:L_{G}^{p}(\Omega_{T})\to{P}$, where P{P} denotes the space of families of random variables $\{X^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$ indexed by the control, with appropriate integrability conditions.
To provide rigorous topological foundations for the homomorphism $\Phi$, we define the target space as a Banach space of parameterized random variables.
Definition 3.2 (Uniform $L^{p}$-Norm).
Let $\mathbf{X}=\{X^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$ be a family of random variables with $X^{\theta}\in L^{p}(P)$ for each $\theta$. The uniform norm is defined as:
$\|\mathbf{X}\|_{\mathcal{P}_{p}}:=\sup_{\theta\in\mathcal{A}^{\Theta}}\|X^{\theta}\|_{L^{p}(\mathbb{P})}.$
Definition 3.3 (Parameterized Cylinder Functions).
The space of parameterized cylinder functions is:
$\mathcal{C}_{par}=\left\{\left\{\varphi(B_{t_{1}}^{\theta},\ldots,B_{t_{n}}^{\theta})\right\}_{\theta\in\mathcal{A}^{\Theta}}:\varphi\in C_{l.Lip}(\mathbb{R}^{n}),\ t_{i}\geq 0\right\}.$
Definition 3.4 (Target Space $\mathcal{P}_{p}$).
The target space is the completion:
$\mathcal{P}_{p}:=\overline{\mathcal{C}_{par}}^{\|\cdot\|_{\mathcal{P}_{p}}},$
which forms a Banach space under the uniform norm.
Remark 3.1.
With this topology, the convergence condition in Definition 3.4 becomes standard Banach space convergence:
$\|\Phi(X)-\Phi(\xi_{n})\|_{\mathcal{P}_{p}}\to 0,$
ensuring the well-definedness of $\Phi$ on $L_{G}^{p}$.
Definition 3.5 (Definition on Cylinder Functions).
For any cylinder function $\xi=\varphi(B_{t_{1}\wedge T},\ldots,B_{t_{n}\wedge T})\in\mathcal{C}_{T}$, we define:
$\Phi(\xi):=\left\{\xi^{\theta}\right\}_{\theta\in\mathcal{A}^{\Theta}},\quad\text{where}\quad\xi^{\theta}:=\varphi(B_{t_{1}\wedge T}^{\theta},\ldots,B_{t_{n}\wedge T}^{\theta}).$
A crucial property of this definition is the preservation of the G-norm, which follows from the definition of G-expectation as a supremum over controls.
Lemma 3.6 (Isometry on Cylinder Functions).
For any cylinder function $\xi\in\mathcal{C}_{T}$, we have:
$\hat{E}[|\xi|^{p}]=\sup_{\theta\in\mathcal{A}^{\Theta}}E_{P}[|\xi^{\theta}|^{p}].$
Consequently, $\|\xi\|_{L_{G}^{p}}=\sup_{\theta}\|\xi^{\theta}\|_{L^{p}(P)}$.
Proof.
This follows directly from the representation of the G-expectation for cylinder functions. See[1]. ∎
This isometry property allows us to extend $\Phi$ to the entire space $L_{G}^{p}(\Omega_{T})$ by continuity.
Definition 3.7 (Extension by Continuity).
For any $X\in L_{G}^{p}(\Omega_{T})$, let $\{\xi_{n}\}\subset\mathcal{C}_{T}$ be a sequence such that $\|X-\xi_{n}\|_{L_{G}^{p}}\to 0$. We define:
$\Phi(X):=\lim_{n\to\infty}\Phi(\xi_{n}),$
where the limit is taken in the sense that:
$\sup_{\theta\in\mathcal{A}^{\Theta}}\|X^{\theta}-\xi_{n}^{\theta}\|_{L^{p}(P)}\to 0\quad\text{as}\quad n\to\infty.$
3.3 Preservation of Algebraic Operations
We now prove that the homomorphism $\Phi$ preserves the algebraic operations of addition, scalar multiplication, and multiplication.
Theorem 3.8 (Preservation of Addition).
For any $X,Y\in L_{G}^{p}(\Omega_{T})$, we have:
$\Phi(X+Y)=\Phi(X)+\Phi(Y).$
Proof.
Let $\{\xi_{n}\},\{\eta_{n}\}\subset\mathcal{C}_{T}$ be sequences approximating $X$ and $Y$, respectively. Then $\{\xi_{n}+\eta_{n}\}$ is a sequence of cylinder functions approximating $X+Y$.
By the definition on cylinder functions, for each $n$:
$\Phi(\xi_{n}+\eta_{n})=\{(\xi_{n}+\eta_{n})^{\theta}\}=\{\xi_{n}^{\theta}+\eta_{n}^{\theta}\}=\Phi(\xi_{n})+\Phi(\eta_{n}).$
Taking the limit as $n\to\infty$ on both sides, and using the continuity of $\Phi$ (which follows from its isometry property), we obtain:
$\Phi(X+Y)=\lim_{n\to\infty}\Phi(\xi_{n}+\eta_{n})=\lim_{n\to\infty}[\Phi(\xi_{n})+\Phi(\eta_{n})]=\Phi(X)+\Phi(Y).$
Note that the last equality holds because the limit of a sum is the sum of the limits in the topology of uniform convergence in $L^{p}(P)$ over $\theta$. ∎
Theorem 3.9 (Preservation of Scalar Multiplication).
For any $X\in L_{G}^{p}(\Omega_{T})$ and $\lambda\in\mathbb{R}$, we have:
$\Phi(\lambda X)=\lambda\Phi(X).$
Proof.
The proof is analogous. Let $\{\xi_{n}\}\subset\mathcal{C}_{T}$ approximate $X$. Then $\{\lambda\xi_{n}\}$ approximates $\lambda X$. For each $n$:
$\Phi(\lambda\xi_{n})=\{(\lambda\xi_{n})^{\theta}\}=\{\lambda\xi_{n}^{\theta}\}=\lambda\Phi(\xi_{n}).$
Taking limits yields the result. ∎
Theorem 3.10 (Preservation of Multiplication).
For any $X,Y\in L_{G}^{p}(\Omega_{T})$, where $p\geq 2$,we have:
$\Phi(X\cdot Y)=\Phi(X)\cdot\Phi(Y).$
Proof.
Let $\{\xi_{n}\},\{\eta_{n}\}\subset\mathcal{C}_{T}$ approximate $X$ and $Y$, respectively. Note that the product $\xi_{n}\eta_{n}$ is also a cylinder function (by considering a function that multiplies the corresponding components).
By the definition on cylinder functions:
$\Phi(\xi_{n}\eta_{n})=\{(\xi_{n}\eta_{n})^{\theta}\}=\{\xi_{n}^{\theta}\eta_{n}^{\theta}\}=\Phi(\xi_{n})\cdot\Phi(\eta_{n}).$
We now show that $\{\xi_{n}\eta_{n}\}$ approximates $X\cdot Y$ in the parameterized framework. For each $\theta\in\mathcal{A}^{\Theta}$, applying the classical Hölder inequality:
$\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}\eta_{n}^{\theta}-X^{\theta}Y^{\theta}|^{p/2}]\leq\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}|^{p}]^{1/2}\mathbb{E}_{\mathbb{P}}[|\eta_{n}^{\theta}-Y^{\theta}|^{p}]^{1/2}+\mathbb{E}_{\mathbb{P}}[|Y^{\theta}|^{p}]^{1/2}\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}-X^{\theta}|^{p}]^{1/2}.$
Since $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}-X^{\theta}|^{p}]\to 0$ and $\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[|\eta_{n}^{\theta}-Y^{\theta}|^{p}]\to 0$, and the sequences $\{\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}|^{p}]\}$ and $\{\mathbb{E}_{\mathbb{P}}[|Y^{\theta}|^{p}]\}$ are uniformly bounded in $\theta$, we obtain:
$\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}_{\mathbb{P}}[|\xi_{n}^{\theta}\eta_{n}^{\theta}-X^{\theta}Y^{\theta}|^{p/2}]\to 0.$
This uniform convergence in $L^{p/2}(\mathbb{P})$ over $\theta$ implies:
$\Phi(X\cdot Y)=\lim_{n\to\infty}\Phi(\xi_{n}\eta_{n})=\lim_{n\to\infty}[\Phi(\xi_{n})\cdot\Phi(\eta_{n})]=\Phi(X)\cdot\Phi(Y).$
Remark 3.2.
The preservation of continuous function composition $\Phi(\phi(X))=\phi(\Phi(X))$ relies on the existence of a common subsequence $\{\xi_{n_{k}}\}$ of cylinder functions with specific convergence properties. In the original G-expectation space, this subsequence converges quasi-surely to $X$, defining an $L^{p}_{G}$ equivalence class. Crucially, in the parameterized space, for each $\theta\in\mathcal{A}^{\Theta}$, the same subsequence converges almost everywhere to $X^{\theta}$ under the probability measure $\mathbb{P}$, while also maintaining $L^{p}$-norm convergence uniformly in $\theta$.
3.4 Analytic Structure Preservation
The homomorphism also preserves analytic properties:
Theorem 3.11 (Analytic homomorphism).
@@ 3.4 Analytic Structure Preservation
    Conditional Expectation: $\hat{\mathbb{E}}_{t}[X]=\sup\limits_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
@@
The homomorphism extends to differential operations:
@@ 3.4.1 Itô Calculus Preservation
Theorem 3.12 (Itô Calculus homomorphism).
@@
    $\Phi\left(\int_{0}^{t}dB_{s}\right)=\left\{\int_{0}^{t}\theta_{s}dW_{s}\right\}$
@@
3.4.2 Itô Calculus homomorphism (Integral Form)
@@
This confirms the homomorphism preserves the integral structure of Itô calculus in L²-sense. ∎
3.5 Schematic of G-SDE Analysis Steps under the Parametric Framework
@@
Remark 3.3.
In the parametric framework for solving G-SDEs, beyond performing conventional operations, it is crucial to ensure that all convergences hold uniformly in the control parameter with respect to the $L^{p}$-norm. This uniform convergence in $\theta$ is precisely what guarantees the well-definedness of the inverse mapping $\Phi^{-1}$, allowing the reconstruction of a single $G$-random variable from the parameterized family $\{X_{t}^{\theta}\}_{\theta\in\mathcal{A}^{\Theta}}$.
Essentially, our approach consolidates the analytical complexity solely into the final stage of uniform estimation, rather than dispersing it throughout the computational process. These uniformity conditions are not only necessary but also economically minimal—they represent the least restrictive requirements to preserve the structural integrity of the homomorphism. This underscores the operational optimality of our parametric representation: it leverages classical tools to their fullest extent while seamlessly bridging to the $G$-expectation framework through a single, coherent consistency analysis.
By adhering to this principle, the parametric representation maintains its mathematical meaning and provides a robust foundation for aggregating the solution family into a universal process on the path space.
@@ 3.6 Algebraic and Analytic Structure Preservation
The homomorphism preserves all algebraic and analytic operations:
@@
The sublinear expectation operators are related through the injective homomorphism as follows:
@@
$\hat{\mathbb{E}}[X]$ $\sup\limits_{\theta\in\mathcal{A}_{0,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}[X^{\theta}]$
$\hat{\mathbb{E}}_{t}[X]$ $\sup\limits_{\theta\in\mathcal{A}_{t,T}^{\Theta}}\mathbb{E}_{\mathbb{P}}[X^{\theta}\mid\mathcal{F}_{t}]$
Note that the control sets differ between the unconditional and conditional expectations: the unconditional expectation uses controls over the entire interval $[0,T]$, while the conditional expectation uses controls only over the remaining interval $[t,T]$.
@@
    Analytic Continuity: The homomorphism preserves convergence properties:
@@ 4 How to Solve G-Stochastic Differential Equations
$dX_{t}^{\theta}=b(t,X_{t}^{\theta})dt+h(t,X_{t}^{\theta})\theta_{t}^{2}dt+\sigma(t,X_{t}^{\theta})\theta_{t}dW_{t}$ (4.1)
@@
$dX_{t}=b(t,X_{t})dt+h(t,X_{t})d\langle B\rangle_{t}+\sigma(t,X_{t})dB_{t}$ (4.2)
where $B_{t}$ is a G-Brownian motion.
4.1 Existence and Uniqueness for the Parameterized Family
Theorem 4.1 (Classical SDE Solution with Quadratic Variation Term).
@@ 4.1 Existence and Uniqueness for the Parameterized Family
Then for each $\theta\in\mathcal{A}^{\Theta}$, there exists a unique strong solution $X_{t}^{\theta}$ to (4.1), and the solution family satisfies the uniform $L^{2}$-estimate:
$\sup_{\theta\in\mathcal{A}^{\Theta}}\mathbb{E}\left[\sup_{0\leq t\leq T}|X_{t}^{\theta}|^{2}\right]\leq C(1+|X_{0}|^{2})e^{CT}$ (4.6)
for some constant $C>0$ depending only on $K$ and $M$.
@@
We prove this via Picard iteration with uniform estimates. Define the Picard iteration scheme:
$X_{t}^{\theta,(0)}=X_{0},\quad X_{t}^{\theta,(n+1)}=X_{0}+\int_{0}^{t}b(s,X_{s}^{\theta,(n)})ds+\int_{0}^{t}h(s,X_{s}^{\theta,(n)})\theta_{s}^{2}ds+\int_{0}^{t}\sigma(s,X_{s}^{\theta,(n)})\theta_{s}dW_{s}.$
Step 1: Uniform $L^{2}$-bounds for iterates. For fixed $\theta$ and $n$, by Itô’s isometry and the growth conditions:
$\mathbb{E}[|X_{t}^{\theta,(n+1)}|^{2}]$ $\leq 4|X_{0}|^{2}+4\mathbb{E}\left[\left|\int_{0}^{t}b(s,X_{s}^{\theta,(n)})ds\right|^{2}\right]+4\mathbb{E}\left[\left|\int_{0}^{t}h(s,X_{s}^{\theta,(n)})\theta_{s}^{2}ds\right|^{2}\right]$
$\quad+4\mathbb{E}\left[\left|\int_{0}^{t}\sigma(s,X_{s}^{\theta,(n)})\theta_{s}dW_{s}\right|^{2}\right]$
$\leq 4|X_{0}|^{2}+4t\mathbb{E}\left[\int_{0}^{t}|b(s,X_{s}^{\theta,(n)})|^{2}ds\right]+4t\mathbb{E}\left[\int_{0}^{t}|h(s,X_{s}^{\theta,(n)})\theta_{s}^{2}|^{2}ds\right]$
$\quad+4\mathbb{E}\left[\int_{0}^{t}|\sigma(s,X_{s}^{\theta,(n)})\theta_{s}|^{2}ds\right]$
$\leq 4|X_{0}|^{2}+4tK^{2}\mathbb{E}\left[\int_{0}^{t}(1+|X_{s}^{\theta,(n)}|)^{2}ds\right]+4tM^{4}K^{2}\mathbb{E}\left[\int_{0}^{t}(1+|X_{s}^{\theta,(n)}|)^{2}ds\right]$
$\quad+4M^{2}K^{2}\mathbb{E}\left[\int_{0}^{t}(1+|X_{s}^{\theta,(n)}|)^{2}ds\right]$
$=4|X_{0}|^{2}+4\left[t(1+M^{4})K^{2}+M^{2}K^{2}\right]\mathbb{E}\left[\int_{0}^{t}(1+|X_{s}^{\theta,(n)}|)^{2}ds\right]$
$\leq 4|X_{0}|^{2}+8\left[T(1+M^{4})K^{2}+M^{2}K^{2}\right]\left(t+\int_{0}^{t}\mathbb{E}[|X_{s}^{\theta,(n)}|^{2}]ds\right).$
Let $C_{1}=8\left[T(1+M^{4})K^{2}+M^{2}K^{2}\right]$. Then:
$\mathbb{E}[|X_{t}^{\theta,(n+1)}|^{2}]\leq 4|X_{0}|^{2}+C_{1}\left(t+\int_{0}^{t}\mathbb{E}[|X_{s}^{\theta,(n)}|^{2}]ds\right).$
By Gronwall’s inequality, there exists $C_{2}>0$ such that:
$\sup_{n\geq 0}\mathbb{E}[|X_{t}^{\theta,(n)}|^{2}]\leq C_{2}(1+|X_{0}|^{2})e^{C_{2}t}.$
Step 2: Uniform convergence of iterates. For $n\geq 1$, consider:
$\mathbb{E}[|X_{t}^{\theta,(n+1)}-X_{t}^{\theta,(n)}|^{2}]$
$\leq 3\mathbb{E}\left[\left|\int_{0}^{t}(b(s,X_{s}^{\theta,(n)})-b(s,X_{s}^{\theta,(n-1)}))ds\right|^{2}\right]$
$\quad+3\mathbb{E}\left[\left|\int_{0}^{t}(h(s,X_{s}^{\theta,(n)})-h(s,X_{s}^{\theta,(n-1)}))\theta_{s}^{2}ds\right|^{2}\right]$
$\quad+3\mathbb{E}\left[\left|\int_{0}^{t}(\sigma(s,X_{s}^{\theta,(n)})-\sigma(s,X_{s}^{\theta,(n-1)}))\theta_{s}dW_{s}\right|^{2}\right]$
$\leq 3\left[t+tM^{4}+M^{2}\right]K^{2}\int_{0}^{t}\mathbb{E}[|X_{s}^{\theta,(n)}-X_{s}^{\theta,(n-1)}|^{2}]ds.$
Let $C_{3}=3(T+TM^{4}+M^{2})K^{2}$ and define $\phi_{n}(t)=\mathbb{E}[|X_{t}^{\theta,(n+1)}-X_{t}^{\theta,(n)}|^{2}]$. Then:
$\phi_{n}(t)\leq C_{3}\int_{0}^{t}\phi_{n-1}(s)ds.$
By induction, $\phi_{n}(t)\leq\frac{(C_{3}t)^{n}}{n!}\phi_{0}(T)$. Since $\sum_{n=0}^{\infty}\sqrt{\frac{(C_{3}T)^{n}}{n!}}<\infty$, the sequence $\{X_{t}^{\theta,(n)}\}$ converges uniformly in $L^{2}$ to some limit $X_{t}^{\theta}$.
Step 3: Verification that limit solves SDE. The limit $X_{t}^{\theta}$ satisfies:
$X_{t}^{\theta}=X_{0}+\int_{0}^{t}b(s,X_{s}^{\theta})ds+\int_{0}^{t}h(s,X_{s}^{\theta})\theta_{s}^{2}ds+\int_{0}^{t}\sigma(s,X_{s}^{\theta})\theta_{s}dW_{s}\quad\text{in }L^{2}.$
The stochastic integral is well-defined since:
$\mathbb{E}\left[\int_{0}^{T}|\sigma(s,X_{s}^{\theta})\theta_{s}|^{2}ds\right]\leq M^{2}K^{2}\mathbb{E}\left[\int_{0}^{T}(1+|X_{s}^{\theta}|)^{2}ds\right]<\infty.$
Step 4: Uniform $L^{2}$-estimate for solutions. Taking supremum and expectation in the SDE:
$\mathbb{E}\left[\sup_{0\leq t\leq T}|X_{t}^{\theta}|^{2}\right]$ $\leq 4|X_{0}|^{2}+4\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\int_{0}^{t}b(s,X_{s}^{\theta})ds\right|^{2}\right]$
$\quad+4\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\int_{0}^{t}h(s,X_{s}^{\theta})\theta_{s}^{2}ds\right|^{2}\right]$
$\quad+4\mathbb{E}\left[\sup_{0\leq t\leq T}\left|\int_{0}^{t}\sigma(s,X_{s}^{\theta})\theta_{s}dW_{s}\right|^{2}\right]$
$\leq 4|X_{0}|^{2}+4T\mathbb{E}\left[\int_{0}^{T}|b(s,X_{s}^{\theta})|^{2}ds\right]+4T\mathbb{E}\left[\int_{0}^{T}|h(s,X_{s}^{\theta})\theta_{s}^{2}|^{2}ds\right]$
$\quad+16\mathbb{E}\left[\int_{0}^{T}|\sigma(s,X_{s}^{\theta})\theta_{s}|^{2}ds\right]$
$\leq 4|X_{0}|^{2}+4TK^{2}\mathbb{E}\left[\int_{0}^{T}(1+|X_{s}^{\theta}|)^{2}ds\right]+4TM^{4}K^{2}\mathbb{E}\left[\int_{0}^{T}(1+|X_{s}^{\theta}|)^{2}ds\right]$
$\quad+16M^{2}K^{2}\mathbb{E}\left[\int_{0}^{T}(1+|X_{s}^{\theta}|)^{2}ds\right]$
$=4|X_{0}|^{2}+4\left[T(1+M^{4})K^{2}+4M^{2}K^{2}\right]\mathbb{E}\left[\int_{0}^{T}(1+|X_{s}^{\theta}|)^{2}ds\right]$
$\leq 4|X_{0}|^{2}+8\left[T(1+M^{4})K^{2}+4M^{2}K^{2}\right]\left(T+\int_{0}^{T}\mathbb{E}[|X_{s}^{\theta}|^{2}]ds\right).$
By Gronwall’s inequality, there exists $C>0$ such that:
$\mathbb{E}\left[\sup_{0\leq t\leq T}|X_{t}^{\theta}|^{2}\right]\leq C(1+|X_{0}|^{2})e^{CT}.$
Crucially, the constant $C$ depends only on $K$, $M$, and $T$, not on the specific control $\theta$.
Step 5: Uniqueness. If $X_{t}^{\theta}$ and $Y_{t}^{\theta}$ are two solutions, then:
$\mathbb{E}[|X_{t}^{\theta}-Y_{t}^{\theta}|^{2}]\leq 3(T+TM^{4}+M^{2})K^{2}\int_{0}^{t}\mathbb{E}[|X_{s}^{\theta}-Y_{s}^{\theta}|^{2}]ds.$
By Gronwall’s inequality, $\mathbb{E}[|X_{t}^{\theta}-Y_{t}^{\theta}|^{2}]=0$ for all $t$. ∎
Theorem 4.2 (G-SDE Solution via Parameterization).
@@
1. For each $\theta\in\mathcal{A}^{\Theta}$, solve the classical SDE (4.1)
2. The boundedness of $\Theta$ ensures uniform estimates: